Nordic Optical Telescope

Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook

Follow us on Twitter



next up previous contents
Next: Instruments Up: AiC Report to NOT Previous: Down Time

Telescope

Pointing

Telescope pointing is now at 8.42'' rms. New automatic pointing model data collection routine installed.

Mirror cleaning

Combined M1, M2 reflectivity was measured at 60% in January 2001 down from 66% in May 2000. Given the demands on technical time to be made by commissioning of NOTCam and MOSCA, it was decided not to realuminize M1 this summer, but we will explore the possibility of realuminizing M2 as time allows. The reflectivity testing scheme is under suspicion as indications of an uneven decay of M2 reflectivity have not been borne out by strongly defocussed stellar images which appear quite uniform. (PB)gif

Mirror Damping

Experiments using both the experimental adaptive optics system JOSE, and raw data from the telescope load cells indicate the presence of a vibration mode in the telescope system of around 12 Hz (probably generated by the active optics system). A DSP-based digital filter system will be implemented in the previously abandoned load-cell/voice-coil mirror damping loop which will have a natural frequency in this area. The visiting Swedish electronics student Anders Magnusson will assist in this project. (GC)

See Appendix A for details.

TCS II

Some progress has been made with the TCS upgrade project. Hardware I/O tests are complete and coding has begun. Delays due to normal TCS work now imply a project completion date of July 2003. (IS)

TMS

Project on hold due to lack of manpower.

Building crashes

The bursts of building crashes experienced in December 2000 and February 2001 have been cause for concern. Their definitive cause is still not known, but a number of possible causes have been addressed already.

The December instances appeared to have been associated with noise in the building/telescope position sensors, as a small modification to the electrical grounding of these devices seemed to have a positive effect. One curious characteristic of these crashes was that the building and telescope remained in alignment throughout.

The February crashes occurred after the ( tex2html_wrap_inline394 200kg) thermal insulation jacket was installed on FIES, and were more normal in that building and telescope were found to be out of alignment after each crash. Investigations showed that the hydrostatic bearing pads were out of alignment, one seriously so. This was probably at least contributory as the azimuth currents during slews were found to be anomalously high indicating excessive friction. Discussions with Torben Andersen and Niels Christian-Jessen helped us with diagnosis and brought to light the original alignment jigs. Two of the three the bearings were then realigned (the third is not perfect, but within spec) and appear to be stable. Azimuth currents are now more normal.

Occasional unexplained building crashes (without telescope/building misalignment) suggest that all is still not well and investigations continue. The latest theory is that they are not building crashes at all, but arise from anomolous responses from the safety system, causing immediate power downs that appear like building crashes. As it is, the crashes all seem to take place without the building and telescope losing alignment as they would in the case of a true crash.

FIES, as a possible cause, was removed from its mount during this work and remains off until we can be sure we have solved the problem and may therefore monitor its effects more confidently. Expert advice has been sought on the ability of the hydrostatic bearing to carry its additional load and appears optimistic at the time of writing.


next up previous contents
Next: Instruments Up: AiC Report to NOT Previous: Down Time

Tim Abbott, AiC
Wed Mar 21 12:56:05 GMT 2001

Back to top Last modified: 21-Mar-2001